OpenStudy (loser66):

How is it? \[\sum_{n\geq 0}(\dfrac{1}{n+1})^2 =\sum_{n\geq 1}\dfrac{1}{n^2}\] Please, help

OpenStudy (loser66):

@reemii

OpenStudy (reemii):

If you write the first terms of both series, you will see something...

OpenStudy (reemii):

Other method (more formal than just "seeing"): In the LHS, set \(k=n+1\) and see how the expression becomes in terms of \(k\).

OpenStudy (loser66):

Yeah.... I "see" it now. Thank you so much. One more question

OpenStudy (loser66):

why \[\int_0^1\int_0^1 \dfrac{1}{1-xy}dxdy\] is a square? 

OpenStudy (reemii):

The integral is not a square. Computing the integral gives a value. It is the "domain of integration" that is a square. The domain of integration is described at the limits of the integral symbols \(\int\).

OpenStudy (loser66):

but when x=1,y=1, the function is undefined, right?

OpenStudy (reemii):

It is possible that the value of the integral is \(+\infty\). Yes, the function can be undefined at the limits, but that means that you have to compute a "generalized integral". Just as in this case: \[ \int_0^1 \frac1{\sqrt x}\, dx \] Not defined at x=0 but the integral has a finite value.

OpenStudy (loser66):

I read three time pi^2/6 with its proof. I don't get why they have to use the transformation to calculate \(\zeta (2)\). Can you explain me?

OpenStudy (reemii):

I can only see that they succeed in computing it.. How would you evaluate the integral?

OpenStudy (loser66):

I understand all of the proof, just want to make clear why they use u,v plane to calculate the integral instead of x,y plane.

OpenStudy (loser66):

Is it not that because the function is undefined when x=y=1. (As they states, it is a hole there)?

OpenStudy (reemii):

They make a change of variables -you might choose another if you want- that leads to a cool expression.. and for which the integral can be computed. Maybe if you're lucky you can find another change of variable that works fine.

OpenStudy (loser66):

hehehe.. NO, I am not that good, I just want to completely understand it.

OpenStudy (loser66):

After transfer to uv plane, the function becomes \(\dfrac{1}{1-u^2+v^2}\) Hence when \(u^2-v^2=1\) , it is undefined, but at that time, the function is out of the new domain integration, right?

OpenStudy (loser66):

No, it is not, when u =1, v =0 , we have u^2-v^2 =1, still undefined.

OpenStudy (loser66):

That is why I want to know the reason they have to transform to u, v plane.

OpenStudy (reemii):

:-o laptop crash.. Actually, you're right to be careful about the domain of the function. Here, it is an improper integral. https://en.wikipedia.org/wiki/Improper_integral The integral still makes sense. The change of variable doesn't change the property of the integral being improper. It just changes the expression of the function to integrate. (the limits too change...) Since the authors are professional mathematicians, they do not take the time to make the details about improper integrals that students would make. (computing limits)

OpenStudy (loser66):

Thanks for the link. I do appreciate. Let me read and digest it first. :)